OpenStudy (bowtie_man123):

Please Help WILL MEDDLE AND FAN 2/3m=-8

OpenStudy (anonymous):

you do the recipicol on both sides of the equation, cross cancel then multiply. 3/2 x 2/3m=-8/1 x 3/2 m = -8/1 x 3/2 m=-12/1 hope that made since.

OpenStudy (bowtie_man123):

yes thank you very much

OpenStudy (bowtie_man123):

i melded and fanned you